# Principal Component Flow Map Learning of PDEs from incomplete, limited, and noisy data

### v1

@@ 4.2 Example 2: 1D wave system with missing variables
Validation testing also uses $100$ trajectories drawn similarly to (40). On the left and center of Figures 8 (noiseless) and 9 (noisy), an example test trajectory is shown at various prediction times, and on the right the average absolute $\ell_{2}$ error the test trajectories is shown over $150$ and $100$ time steps, respectively. We see that the PC-FML strategies stay relatively stable and smooth over the prediction outlook, while the nodal FML deteriorates almost immediately. Particularly in the noisy case, deviations here are similar to those observed in numerical methods, whereby small errors in initial conditions accumulate and while the solution exhibits “wave-like” behavior, it is “displaced” from the appropriate trajectory. Additionally, the denoising behavior of PC-FML approach is evident.
@@
Figure 8: Ex. 2 (noiseless): Example trajectory (left, center) and average error (right).
Figure 9: Ex. 2 (noisy): Example trajectory (left, center) and average error (right).
@@ 4.3 Example 3: 2D wave equation on non-uniform grid with missing information
Finally, we consider the two-dimensional wave equation
@@
To generate training data, first we solve the PDE until $T=4$ ($400$ time steps) from $N_{traj}=100$ randomized initial conditions given by
@@
Figure 10: Ex. 3: Non-uniform measurement grid.
@@
Validation testing also uses $100$ trajectories drawn similarly to (42). On the left and center of Figures 11 (noiseless) and 12 (noisy), an example test trajectory is shown at various prediction times, and on the right the average absolute $\ell_{2}$ error the test trajectories is shown over $1000$ and $200$ time steps, respectively. We see that the PC-FML strategies stay relatively stable and smooth over the prediction outlook, while the nodal FML noisily deteriorates almost immediately and never actually exhibits “wave-like” behavior. Additionally, the denoising behavior of PC-FML approach is evident. Generally, the results show that despite its simplicity, the SVD-motivated PC-FML approach yields accurate approximation to nonlinear data.
@@
Figure 11: Ex. 3 (noiseless): Example trajectory (left, center) and average error (right).
Figure 12: Ex. 3 (noisy): Example trajectory (left, center) and average error (right).
@@ 5 Conclusion
We have presented a computational technique in the Flow Map Learning (FML) family for modeling the evolution of unknown PDEs from limited, noisy, and incomplete solution data using a novel NN architecture with significantly reduced parameterization, notably enabling hundredfold lower training data requirements. In particular, the architecture achieves this drastic decrease in parameters by first learning a common reduced linear approximation to reduce high-dimensional spatial measurements to a few important components, and then conducting the dynamics learning in this reduced basis which requires a lower complexity network due to its ODE form. Future work will focus on efficient learning nonlinear reduced representations in the context of data-driven FML of PDEs.

### v2

@@ 4.3 Example 3: 1D wave system with missing variables
Validation testing also uses $100$ trajectories drawn similarly to (40). On the left and center of Figures 10 (noiseless) and 11 (noisy), an example test trajectory is shown at various prediction times, and on the right the average absolute $\ell_{2}$ error the test trajectories is shown over $300$ time steps. We see that the PC-FML strategies stay relatively stable and smooth over the prediction outlook, while the nodal FML deteriorates almost immediately. Particularly in the noisy case, deviations here are similar to those observed in numerical methods, whereby small errors in initial conditions accumulate and while the solution exhibits “wave-like” behavior, it is “displaced” from the appropriate trajectory. Additionally, the denoising behavior of PC-FML approach is evident.
@@
Figure 10: Ex. 3 (noiseless 1D wave): Example trajectory (left, center) and average error (right).
Figure 11: Ex. 3 (noisy 1D wave): Example trajectory (left, center) and average error (right).
4.4 Example 4: 1D shallow water system with missing variables
Next we consider the nonlinear shallow water equations
$\begin{split}h_{t}&+(hu)_{x}=0\\ (hu)_{t}&+\left(hu^{2}+\frac{9.8}{2}h^{2}\right)_{x}=0\end{split}$ (41)
where $h$ and $u$ represent height and velocity of a fluid, respectively. We operate on the domain $\Omega=[0,10]$ with periodic boundary conditions. We consider the case when only data on the solution component $h$ are available (i.e. $N_{obs}=1$) and aim at constructing an accurate NN predictive model for the evolution of $h$.
To generate training data, first we numerically solve the PDE until $T=5$ ($626$ time steps) from $N_{traj}=100$ randomized initial conditions, representing a sinusoidal perturbation in the fluid height with zero initial velocity, given by
$\begin{split}h(x,0)&=1+\alpha\sin\left(\frac{2\pi x}{10}\right),\quad u(x,0)=0\end{split}$ (42)
with $\alpha\sim U[0.1,1.0]$. Then, one random chunk of length $N_{mem}+N_{rec}$ per trajectory from $h$ only are recorded into our training data. The solutions are observed on a uniform grid of $N_{grid}=200$ points.
The method from Section 3.4 is used to determine an appropriate reduced basis of dimension $N_{red}=52$ for the noiseless case and $N_{red}=6$ noisy case of $\sigma=10^{-2}$. Figure 3 shows the details. Particularly in this example, the dimension $N_{red}$ required to approximate solution data with high accuracy is quite large, resulting in $P_{in}$ and $P_{out}$ operators of size $200\times 52$ and $52\times 200$ and presenting an input to the flow map network $\mathbf{M}$ of length $1040$. Because of these dimensionality challenges and our focus on rapid and inexpensive local simulation of scientific data, in this case we consider only the Fixed approach, thereby reducing the training burden on PC-FML.
Validation testing also uses $100$ trajectories drawn similarly to (42). On the left and center of Figures 12 (noiseless) and 13 (noisy), an example test trajectory is shown at various prediction times, and on the right the average absolute $\ell_{2}$ error the test trajectories is shown over $100$ and $275$ time steps, respectively. In the noiseless case, we see that the nodal FML model fails to evolve, while the Fixed PC-FML model follows the correct evolution before being infected with small perturbations in the prediction. In the noisy case, the Fixed PC-FML performs even better, with denoised predictions with some accuracy up until $t=2.2$. Since the shallow water equations represent a hyperbolic conservation law, we expect and plan that improvements to these results can be made by integrating into PC-FML techniques from [6, 20] that enforce conservation.
Table 1 shows that the nodal FML model uses between $10$ and $25$ times more parameters than the PC-FML approach.
Figure 12: Ex. 4 (noiseless 1D shallow water): Example trajectory (left, center) and average error (right).
Figure 13: Ex. 4 (noisy 1D shallow water): Example trajectory (left, center) and average error (right).
@@ 4.5 Example 5: 2D wave equation on non-uniform grid with missing information
In our first two-dimensional example, we consider the wave equation
@@
To generate training data, first we numerically solve the PDE until $T=4$ ($400$ time steps) from $N_{traj}=100$ randomized initial conditions given by
@@
Figure 14: Ex. 5 (2D wave): Non-uniform measurement grid.
@@
Validation testing also uses $100$ trajectories drawn similarly to (44). On the left and center of Figures 15 (noiseless) and 16 (noisy), an example test trajectory is shown at various prediction times, and on the right the average absolute $\ell_{2}$ error the test trajectories is shown over $1000$ and $200$ time steps, respectively. We see that the PC-FML strategies stay relatively stable and smooth over the prediction outlook, while the nodal FML noisily deteriorates almost immediately and never actually exhibits “wave-like” behavior. Additionally, the denoising behavior of PC-FML approach is evident.
@@
Figure 15: Ex. 5 (noiseless 2D wave): Example trajectory (left, center) and average error (right).
Figure 16: Ex. 5 (noisy 2D wave): Example trajectory (left, center) and average error (right).
4.6 Example 6: 2D Navier-Stokes equation
Finally, we consider the nonlinear Navier-Stokes equations in their vorticity-streamfunction formulation
$\begin{split}\omega_{t}+\psi_{y}\omega_{x}-\psi_{x}\omega_{y}&=\nabla^{2}\omega\\ \nabla^{2}\psi+\omega&=0\end{split}$ (45)
where $\omega$ represents vorticity and $\psi$ the streamfunction. We operate on the domain $\Omega=[0,10]^{2}$ with periodic boundary conditions in both directions. Since $\psi$ is directly accessible by solving a Laplace equation if $\omega$ is known, observing only $\omega$ represents a fully-observed system. We aim at constructing an accurate NN predictive model for the evolution of $\omega$, demonstrating the value of PC-FML in the fully-observed case as well.
To generate training data, first we numerically solve the PDE until $T=5$ ($500$ time steps) from $N_{traj}=100$ randomized initial conditions given by
$\begin{split}\omega(x,0)&=e^{-\frac{1}{2}\left((x-x_{0})^{2}+(y-y_{0})^{2}\right)}\end{split}$ (46)
with $x_{0},y_{0}\sim U[2,8]^{2}$. Then, one random chunk of length $N_{mem}+N_{rec}$ per trajectory from $u$ are recorded into the training data. The solutions are observed on a uniform grid of $N_{grid}=50^{2}$ points.
The method from Section 3.4 is used to determine an appropriate reduced basis of dimension $N_{red}=46$ for the noiseless case $(\sigma=0)$ and $N_{red}=26$ for the noisy case $(\sigma=10^{-2})$. Figure 3 shows the details. In particular, we see that the noiseless singular values of the training data drop off gradually after $N_{red}=46$, while noise is dominant after $N_{red}=26$ in the noisy observations. Unlike the linear examples, indicated by the training data generation, there are not obvious optimal basis functions for the initial conditions. Like in Ex. 4, we focus on the Fixed PC-FML approach due to the high dimensionality even in the reduced basis.
Validation testing also uses $100$ trajectories drawn similarly to (46). On the left and center of Figures 17 (noiseless) and 18 (noisy), an example test trajectory is shown at various prediction times, and on the (bottom) right the average absolute $\ell_{2}$ error the test trajectories is shown over $500$ time steps. Fixed PC-FML stays stable and smooth over the prediction outlook, while the nodal FML noisily deteriorates almost immediately, with the comparison particularly stark in the noisy case. Generally, the results show that despite its simplicity, the SVD-motivated PC-FML approach yields accurate approximation to data resulting from a nonlinear PDE.
Table 1 shows for the noiseless case, the nodal FML model uses $12087766$ parameters, while the Fixed PC-FML models use just $53496$. In the noisy case, the nodal FML model uses $6837666$ parameters, while the Fixed PC-FML models use just $18296$, representing a compression of over $300$ times.
Figure 17: Ex. 6 (noiseless Navier-Stokes): Example vorticity fields at $t=2.5$ and average error (bottom right).
Figure 18: Ex. 6 (noisy Navier-Stokes): Example vorticity fields at $t=5$ (center, left) and average error (right).
@@ 5 Conclusion
We have presented a computational technique in the Flow Map Learning (FML) family for modeling the evolution of unknown PDEs from limited, noisy, and incomplete solution data using a novel NN architecture with significantly reduced parameterization, notably enabling hundredfold lower training data requirements and the potential for rapid local simulation of high-dimensional systems. In particular, the architecture achieves this drastic decrease in parameters by first learning a common reduced linear approximation to reduce high-dimensional spatial measurements to a few important components, and then conducting the dynamics learning in this reduced basis which requires a lower complexity network due to its ODE form. Future work will focus on incorporating physics-based constraints to further reduce representations and dynamics in the context of data-driven FML of PDEs.